# Concept Formation and Alignment in Language Models: Bridging Statistical Patterns in Latent Space to Concept Taxonomy

### v1

@@ 4.1 LM-LM alignment
Case Sensitivity Analysis. The T5 model’s vocabulary preserves case information, enabling us to examine how formed concepts align with capitalization differences. We identified 4,328 tokens with varying case appearances (total of 8,887 tokens). We found that for highly granular concepts (k=6), 80% of these tokens belong to the same community. This ratio increases to 85% for k=25 before plateauing, suggesting that case variations generally do not significantly alter the semantic grouping of tokens. This finding supports the notion that the model learns to associate words with their meanings regardless of capitalization, particularly for more abstract or broader concepts (larger k values). However, the initial increase in alignment ratio with increasing k implies that case sensitivity might still play a minor role in differentiating highly specific or nuanced concepts.
4.2 Human-LM alignment
In the previous section, we established the concepts formed in LM models have a degree of alignment with respect to each other. To investigate human-LM alignment, in this section, we focus our analysis on the ALBERT model. While our empirical results are derived from ALBERT, the principles discussed are generalizable to other LLMs, given the LM-LM alignment established earlier (we include human-LM alignment evaluations for GloVe and T5 in Appendix B).
Concept Hierarchy Formation. When examining the broader perspective (i.e., $k$=100), the model primarily found groups of named entities (mostly names of people and places), adverbs, sub-words, some number symbols, and a super-community for ‘other vocabulary’ (appendix C, Figure 4 shows the overall concept hierarchies of the ALBERT vocab). Zooming in further (e.g., $k$=75), these communities revealed more specific clusters that are relatable to the real world. For instance, within named entities, clusters formed for personal vs. location names, even further pinpointing locations by country. Similarly, within the aforementioned super-community, sub-clusters emerged that mapped to various lifestyle aspects, including relationships, social structures, multimedia, and motile creatures. As the granularity level increases (approaching a smaller $k$ value), clusters exhibit a stronger association with word forms.
Concept Alignments. To investigate the presence and extent of human-LM concept alignment, we systematically examine the model’s representations of distinct facets of the external world:
  • Numerical Concepts. This category focuses on numerical values, facilitating a targeted analysis of intra-category relationships, such as value-based ordering.
  • Named Entities (Names and Locations). Primarily comprising proper nouns, this category leverages readily available labeled datasets to validate alignment between the LM’s conceptual groupings and established knowledge bases. It also allows us to investigate if associative relations (e.g., country-nationality) influence community formation.
  • Social Structures. This category reflects the LM’s approach to forming categories that encompass both abstract and concrete nouns, including roles, titles, and processes. This analysis can elucidate its ability to grasp and align complex relationships and concepts.
By analyzing these diverse facets, we aim to comprehensively evaluate the degree to which the LM’s internal representations align with human understanding of the external world.
Table 3: Orderliness score of different communities of the numbers. Years refer to the cluster of numbers between 1816 - 2021. Support refers to the number of samples within the cluster.
Category Top-1 Top-3 Top-5 Support
0-100 0.68 0.89 0.92 100
100-262 0.75 0.94 1.0 115
263-300 0.46 0.78 0.89 37
300-400 0.42 0.71 0.84 82
400+ 0.53 0.75 0.83 110
Years 0.86 0.96 0.98 203
4.2.1 Symbols-Numbers
At a high level, numbers and symbols are completely separated from the rest of the tokens in the embedding space. within themselves they form communities according to (i) years, (ii) integer values, (iii) tokens indicating monetary values (e.g., \$1), (iv) ratio/time (e.g., 2:1, 3:30), (v) fractions (e.g., ’/8’), (vi) large values (e.g., ‘100,000’), and (vii) percentages (e.g., $42\%$). Within these communities, integers form communities based on their hundreds. Within the embedding manifold, we analyzed the spatial distribution of number sequences, particularly focusing on their sequencing. Since the embedding manifold is not guaranteed to be a Euclidean structure [24, 8, 7], conventional distance-based measures of order were not applicable. We define a point $x$ as "locally ordered" on a manifold if it resides within the intersection of the $k$-nearest neighbor sets of both its immediate neighbors, $x+1$ and $x-1$. This localized approach captures the intuitive notion of orderliness within a small neighborhood on the manifold, rather than relying on a global ranking.
Orderliness Score. Table 3 presents the orderliness scores of the communities identified by our algorithm. The "top-1" score represents strict orderliness, where $x+1$ lies exactly between $x$ and $x+2$ in the embedding space. The "top-3" and "top-5" scores measure relaxed orderliness, where $x+1$ falls within the 3 or 5 nearest neighbors of both $x$ and $x+2$, respectively. As the results indicate, all detected communities exhibit a higher degree of local orderliness, regardless of the chosen level of strictness. This finding is significant because it suggests that language models may possess not only the ability to categorize heterogeneous input entities but also the capacity to construct meaning within smaller, internally consistent structures (internally homogeneous sub-structures). This implies the potential for LMs to move beyond simply classifying information to actively interpreting and generating meaning within specific contexts.
Figure 2: Visualization of the identified name and location communities of size larger than 10 entities. UMAP projection as well as kernel density estimate (KDE) plot is used for the visualization.
4.2.2 Names and Locations
@@
To mitigate the subjectivity risk of assessing the semantic structure, we further used external datasets in our evaluations. For Names, we used name-dataset [44] which consists of a comprehensive set of names (730K first names and 983K last names), their associated genders, and their popularity rank for each country. For locations, we used the country-state-city database [11] which contains information on 250 countries, 5K+ states, and 150K+ cities. Table 4(a) shows the high-level communities that our approach detected. Most of the high-level communities are a mix of names/locations associated with certain geographical/cultural regions. Within these clusters, names and locations form distinct subclusters which we discuss in more detail in the following subsections.
Table 4: Evaluation of the location and name tokens for different hierarchy levels. 4(a) presents the precision for high-level communities, while 4(b) and 4(c) show the more granular communities formed within the detected communities in 4(a)
(a) Precision of the largest identified name and location communities with respect to name and location databases. Note that these are at a higher level in the cluster hierarchies. Table 4(b) shows the identified granular sub-clusters and their associated precision.
@@
West-Asia 0.684 390 Human & Location Names
Hispanic/Latino 0.685 282 Human & Location Names
@@
Europe 0.739 215 Human & Location Names
East-Asia 0.741 178 Human & Location Names
(b) The precision of communities based on the identified categories with respect to name-database. First indicate categories of first names last indicates last names
@@
(c) The precision of communities based on the identified categories with respect to the location database.
@@
Names. To determine whether the sub-clusters are associated with names, we pre-filtered the clusters that at least 70% of their tokens are in the top 1000 names (of any country), with gender confidence of above 0.8. We primarily use gender confidence to distinguish between first-name from last name clusters. Then we cross-referenced all identified name communities for all countries in the dataset and assigned the country name with the highest score  [^3] For some clusters we picked the second highest country name if the scores were similar. Due to the cosmopolitan nature of countries like the USA, They tend result in a high score across the board. as the cluster names.
@@
Table 4(b) shows the precision score of the identified granular communities. The first column shows the ratio of the community members that are indeed human names (overall precision), while the second column shows the ratio with respect to specific countries. It should be noted that the country-wise scores of the countries with similar cultures and languages were similar. The reported precision score shows a high degree of categorization based on the country/culture of origin.
It should be noted that we only included clusters with sizes larger than 10 and country-wise precision of larger than 0.5 due to space limitation. The list of identified cluster names goes far beyond the aforementioned table. Clusters such as character names from books, mythology, and car brand names were also identified which were not included due to space limitations.
@@
Locations. It should be noted that, although the dataset is somewhat comprehensive, the LM token space associates only one token to each location, leading to an artificial decrease in precision for multi-token location names. Despite these limitations, we were able to identify communities within the input embedding space that are associated not only with the location category but also with specific regions/countries. The precision numbers in Table 4(c) suggest the model groups the locations with respect to the borders of the countries, which, to a certain degree, implies a subjective perception of geographical knowledge aligned with human perception (since it is evaluated against human annotated dataset). This finding suggests that the language model attempts to construct a human-like mental map  [^4] We refer the definition of the mental map to the [American Psychology Association dictionary] (https://dictionary.apa.org/mental-map). It is defined as “a mental representation of the world or some part of it based on subjective perceptions rather than objective geographical knowledge., wherein it approximates borders and associates nationalities with specific clusters/communities. It should be noted that, due to space constraints, we only present the communities containing more than 10 entities  [^5] The complete set of clusters is included as the supplementary material.. Note that dealing with multi-token location names is more challenging. For instance, "Carolina" was correctly clustered within the United States community by the LM, while our reference dataset misclassified it as a city in Brazil. Additionally, not all location names with English spellings are included in the dataset; for example, ’Wurttemberg’ (or ’Nuremberg’), a region within Germany, is absent, leading to an artificial decrease in precision.
4.2.3 Social Structures
@@
Our methodology identifies a cluster of 903 members with the theme of social structures. In order to have a reference dataset, we annotated the dataset using a combination of GPT-4 and human annotators 666 A human annotator was used alongside GPT-4 to correct GPT-4 misclassifications. Label correction by a human happens faster than the label suggestion task. We estimated our approach is more time/cost-effective while resulting in the same quality annotations.. Since analyzing such concepts can accompany a degree of human subjectivity, we feed the datasets to GPT-4 and ask to identify the theme of each category (as shown in Table 5) for mitigating such uncertainity. In order to reduce subjectivity (we formed the class names based on the GPT-4 recommendation. Then, we asked GPT-4 and human annotator to classify each word with respect to given class names (we added another class named “Other" to avoid forcing the annotators to misclassification).
@@
4.3 Knowledge-Reasoning Separation
Table 6: Huggingface ALBERT base model on GLUE, SuperGLUE, and SQUAD tasks. For the baseline, the model was finetuned without altering the embeddings. For the mid-point, the embedding layer entries are assigned the mid-point embedding of their associated community. We used [41] repo for benchmarking.
mid-point GLUE
Method/Tasks mnli mrpc qnli qqp rte sst stsb wnli
baseline 0.827 0.841 0.902 0.858 0.765 0.915 0.872 0.549
mid-point 0.849 0.865 0.910 0.875 0.783 0.922 0.890 0.563
mid-point SuperGLUE SQUAD SQUAD
Method/Tasks boolq cb copa multirc record wic wsc v1 (f1) v2 (f1)
baseline 0.622 0.512 0.59 0.350 0.586 0.595 0.528 83.72 70.9
mid-point 0.621 0.478 0.55 0.372 0.588 0.626 0.634 84.4 74.9
Now, we are interested to see the extent to which the knowledge learned during pretraining is modular and separable from the reasoning mechanisms employed by the language model in downstream tasks. For example, can the knowledge learned during the pretraining phase be selectively removed or modified without significantly impacting the model’s performance on finetuning? The modularity can also impact the effectiveness of Language Model Inversion [35, 49] techniques, which aim to extract private information such as names or other sensitive details learned during the pre-training777 For example, if knowledge is found to be highly modular, it may be possible to develop targeted interventions that obscure or remove specific sensitive information without significantly impacting the model’s overall performance on downstream tasks..
5 Limitations and Risks
Risks: This work provides detailed information about (1) the formed clusters/concept in the input embedding layer, and (2) the separation of memory from reasoning in ALBERT. As the methodology can also be applied to other models, it can potentially facilitate more advanced adversarial attacks and content manipulation in language models.
6 Conclusion and Future Works
This research focused on concept formation and alignment across different language models (LMs) and between LMs and humans. We propose a concept extraction mechanism that uncovers the emergence of distinct conceptual communities within the entire embedding space. This approach gives a more holistic view of conceptual representation than examining specific samples, as it explores the full landscape of concepts embedded within the models.
We examined GloVe, ALBERT, and T5 LMs and found that all of them form communities that can be aligned with the external world categories/concepts. We further investigated various aspects of the emergent conceptual communities, ranging from their numerical sequencing to the interrelationships and associations found within the blend of concrete and abstract concepts. Across various language models and types of concepts, distinct communities emerge in the embedding space. These communities demonstrate a significant alignment with human-understandable concepts, suggesting that statistical patterns, rather than specific architectures, or conceptual domains, drive concept formation across LMs. However, the quality and alignment of the formed concepts is a function of the architectural choices, training data, and tokenization schemas.
We further observed that concept formation in transformer-based LMs occurs as early as the input embedding space. This suggests that the input embeddings themselves encode substantial semantic information, contrary to traditional views where meaning emerges later. Moreover, the emergent concepts within the input embedding layer exhibit a degree of separation from subsequent transformer layers.
Early concept formation has several implications, opening exciting avenues for future line of works:
  • Knowledge transfer across tasks and models: By identifying and leveraging shared conceptual structures across different tasks and models, early concept formation can enable more efficient transfer learning and reduce the need for extensive retraining on new tasks.
  • Improving model architectures, training, and evaluation: Understanding how concepts emerge early in the model can inform the design of architectures that better capture and represent meaning, leading to improved performance and interpretability. This knowledge can also guide the development of more effective training techniques and evaluation metrics.
  • Mitigating training data information leakages and reducing biases: Early concept formation can be utilized to identify potential biases and information leakages in the input embedding layer. This opens the door to the development of token embedding engineering techniques to mitigate these biases without requiring re-train.
Appendix A Models Under Investigation
A.1 ALBERT
ALBERT (A Lite BERT) [23] is a transformer-based model for language representation learning, designed to be more efficient than its predecessor, BERT [10]. While it shares the same basic architecture as BERT, it incorporates two main key modifications:
Note that reducing the number of parameters and sharing information across layers can force the model to learn more general representations, thus indirectly contributing to better concept formation. datasets. ALBERT is pretrained on English text datasets, namely the English Wikipedia and BookCorpus, using self-supervised learning objectives.
Tokenization. It uses Sentencepiece tokenizer [22] on the uncased corpus with a vocabulary size limit of 30K tokens.
A.2 T5
A.3 GloVe
Tokenization. They used Stanford tokenizer  [^10] https://nlp.stanford.edu/software/tokenizer.shtml, a form of BPE tokenization scheme [47, 12] that constructs unigram counts from a corpus, and optionally thresholds the resulting vocabulary based on total vocabulary size (2.2M 1111112196016 cased tokens most frequent words in the case of GloVe embeddings) or minimum frequency count [40].
Appendix B Human-LM Alignments For Glove, T5
We performed our concept extraction algorithm on Glove, ALBERT, and T5 models. We observed that concept formation happens in all the models, however, the quality of the formed concepts is better in ALBERT model. Glove embeddings contain more than 200K tokens, in order to enable apple-to-apple comparison with ALBERT, we found the intersection of these tokens with the huggingface Albert-base-v2 token set, and, applied our concept extraction algorithm. We the formed clusters have weak correlation with the concepts formed in ALBERT model 121212if we assume Glove-ALBERT mapping exists between two clusters if more than half of their members are equivalent, %69 percent of the Glove clusters have a corresponding ALBERT cluster for K=6. (they similarities are stronger for concrete names/entities).
Tables 7 and 8 show the T5 location and name communities detected by our algorithm. High precision numbers for these cluster indicate clear concept formation. However,since the pretrained HuggingFace T5 uses cased token set for the pretraining, the number of tokens in associated with location and names are much smaller than T5 and Glove.
Tables 9 and 10 show the GloVe location and name communities detected by our algorithm. Although, it shows the categories are formed in GloVe embeddings as well, the numbers suggests the quality of the formed categories have less quality than both ALBERT and T5 counterpart.
@@ Appendix B Human-LM Alignments For Glove, T5
Table 7: Location communities founds in T5 tokens. Note that most of the communities are member of 0_0_5, and 0_0_13 super communities.
@@
Politician Names 1.0 N/A N/A 14 0_0_5_13
@@
Appendix C Extracted Concept Hierarchies

### v3

@@ Appendix G Case Sensitivity Analysis
The T5 model’s vocabulary preserves case information, enabling us to examine how formed concepts align with capitalization differences. We identified 4,328 tokens with varying case appearances (total of 8,887 tokens). We found that for highly granular concepts (k=6), 80% of these tokens belong to the same community. This ratio increases to 85% for k=25 before plateauing, suggesting that case variations generally do not drastically alter the semantic grouping of tokens. This finding supports the notion that the model learns to associate words with their meanings regardless of capitalization, particularly for more abstract or broader concepts (larger k values). However, the initial increase in alignment ratio with increasing k implies that case sensitivity might still play a minor role in differentiating highly specific or nuanced concepts.
Appendix H ALBERT Human-Model Aligment
H.1 Names and Locations
@@
To mitigate the subjectivity risk of assessing the semantic structure, we further used external datasets in our evaluations. For Names, we used name-dataset 43 which consists of a comprehensive set of names (730K first names and 983K last names), their associated genders, and their popularity rank for each country. For locations, we used the country-state-city database 12 which contains information on all countries, 5K+ states, and 150K+ cities. Table 9 shows the high-level communities that our approach detected. Most of the high-level communities are a mix of names/locations associated with certain geographical/cultural regions. Within these clusters, names and locations form distinct sub-communities which we discuss in more detail in the following subsections.
Note that since there’s no one-to-one mapping between names/locations in LLM vocabulary and external datasets (e.g., a name may appear in multiple countries, the external dataset is also a superset), recall is less relevant. We prioritize precision to evaluate accuracy, as it better reflects our ability to identify correct matches.
Table 7: Precision of the largest identified name and location communities with respect to name and location databases. Note that these are at a higher level in the cluster hierarchies. Table 9 shows the identified granular sub-clusters and their associated precision.
@@
West-Asia 0.684 390 Human & Location
Hispanic/Latino 0.685 282 Human & Location
@@
Europe 0.739 215 Human & Location
East-Asia 0.741 178 Human & Location
Table 8: Precision of communities based on the identified categories with respect to name-database. First and last indicates category of first and last names.
@@
Table 9: Precision of communities based on the identified categories with respect to the location database.
@@
Names. To determine whether the sub-clusters are associated with names, we pre-filtered the clusters that at least 70% of their tokens are in the top 1000 names (of any country), with gender confidence of above 0.8. We primarily use gender confidence to distinguish between first-name from last-name clusters. Then we cross-referenced all identified name communities for all countries in the dataset and assigned the country name with the highest score  [^9] For some clusters we picked the second highest country name if the scores were similar. Due to the cosmopolitan nature of countries like the USA, They tend result in a high score across the board. as the cluster names.
@@
Table 9 shows the precision score of the identified granular communities. The first column shows the ratio of the community members that are indeed human names (overall precision), while the second column shows the ratio with respect to specific countries. It should be noted that the country-wise scores of the countries with similar cultures and languages were similar. The reported precision score shows a high degree of categorization based on the country/culture of origin.  [^10] It should be noted that we only included clusters with sizes larger than 10 and country-wise precision of more than 0.5 due to space limitation. The list of identified cluster names goes far beyond the aforementioned table. Clusters such as character names from books, mythology, and car brand names were also identified which were not included due to space limitations.
The formation of these clusters in the input embedding space, particularly those containing ethnic minority names, presents an opportunity for token engineering to mitigate potential ethnicity biases. (As our focus here is on interpretability and conceptual groupings and alignment, we provide an example of such a token engineering approach in Appendix 6 for interested readers.)
@@
Locations. It should be noted that, although the dataset is majorly comprehensive, the LLM token space associates only one token to each location, leading to an artificial decrease in precision for multi-token location names. Despite these limitations, we were able to identify communities within the input embedding space that are associated not only with the location category but also with specific regions/countries. The precision numbers in Table 9 suggest that the model groups the locations with respect to the borders of the countries, which, to a certain degree, implies a subjective perception of geographical knowledge aligned with the external world  [^11] We refer the definition of the mental map to American Psychology Association dictionary. It is defined as “a mental representation of the world or some part of it based on subjective perceptions rather than objective geographical knowledge., wherein it approximates borders and associates nationalities with specific clusters/communities. It should be noted that, due to space constraints, we only present the communities containing more than 10 entities  [^12] The complete set of clusters is included as the supplementary material.. Note that dealing with multi-token location names is more challenging. For instance, "Carolina" was correctly clustered within the United States community by the LLM, while our reference dataset misclassified it as a city in Brazil. Additionally, not all location names with English spellings are included in the dataset; for example, ’Wurttemberg’ (or ’Nuremberg’), a region within Germany, is absent, leading to an artificial decrease in precision.
H.2 Social Structures
@@
Our methodology identifies a cluster of 903 members with the theme of social structures. In order to have a reference dataset, we annotated the dataset using a combination of GPT-4 and human annotators 131313 A human annotator was used alongside GPT-4 to correct GPT-4 misclassifications. Label correction by a human happens faster than the label suggestion task. We estimated our approach is more time/cost-effective while resulting in the same quality annotations.. Given the potential human subjectivity in analyzing these concepts, we ask GPT-4 to identify the theme of each category in the datasets. (as shown in Table 10) for mitigating such uncertainity. In order to reduce subjectivity (we formed the class names based on the GPT-4 recommendation. Then, we asked GPT-4 and human annotator to classify each word with respect to given class names (we added another class named "Other" to avoid forcing the annotators to miscalassify).
@@
Appendix I Human-LM Alignments For Glove, T5, Llama
@@
Table 11: Location communities founds in T5 tokens. Note that most of the communities are member of 0_0_5, and 0_0_13 super communities. SOAM stands for South America.
@@
We performed our concept extraction algorithm on Glove, Albert, and T5 models. We observed that conceptual groupings happens in all the models, however, the quality of the formed concepts is better in Albert model. Glove embeddings contain more than 200K tokens, in order to enable apple-to-apple comparison with Albert, we found the intersection of these tokens with the huggingface Albert-base-v2 token set, and, applied our concept extraction algorithm. We the formed clusters have weak correlation with the concepts formed in Albert model 141414if we assume Glove-Albert mapping exists between two clusters if more than half of their members are equivalent, %69 percent of the Glove clusters have a corresponding Albert cluster for K=6. (they similarities are stronger for concrete names/entities).
Tables 11 and 12 show the T5 location and name communities detected by our algorithm. High precision numbers for these cluster indicate clear conceptual groupings. However,since the pretrained HuggingFace T5 uses cased token set for the pretraining, the number of tokens in associated with location and names are much smaller than Albert and Glove. Table 13 shows the Llama3 name communities detected by our algorithm. Tables 14 and 15 show the GloVe location and name communities detected by our algorithm. Although, it shows the categories are formed in GloVe embeddings as well, the numbers suggests the quality of the formed categories have less quality than both Albert and T5 counterpart.
@@
Politicians 1.0 N/A N/A 14 0_0_5_13
Table 13: Examples of Name Entity clusters with size greater than 25 tokens found in Llama3 input embedding size. Note that we only evaluate against external datasets with English named entities.
Type Precision Support Note Cluster names
female 0.757 393 US/UK 0_0_0_0_0_0
male 0.811 291 US/UK 0_0_0_0_0_2
male 0.6639 171 US/UK 0_0_0_0_0_3
male 0.495 105 Saudi/UAE 0_0_0_0_0_5
male 0.6875 53 US/South Africa 0_0_0_0_0_11
male 0.5588 46 UK/Canada 0_0_0_0_0_12
female 0.571 41 US/UK 0_0_0_0_0_13
male 0.551 37 Mexico 0_0_0_0_0_14
male 0.619 30 Netherlands/US 0_0_0_0_0_16
male 0.666 30 US/UK 0_0_0_0_0_17
male 0.653 27 Biblical 0_0_0_0_0_18
last names 0.724 340 US/UK 0_0_0_0_0_1
last names 0.528 135 Canada/US 0_0_0_0_0_4
last names 0.509 65 UK/US 0_0_0_0_0_8
last names 0.717 39 Mexico/Chile 0_0_0_0_6_0
locations 0.75 485 Mostly American 0_0_0_8_0
@@
Appendix J Extracted Concept Hierarchies